blaze Bledsoe is in a 33% tax bracket. Blaze is considering purchasing some bonds. He can purchase tax-exempt bonds offered by Baptist health system that have an interest (coupon) rate of 6 1/4%. or he can purchase community health systems taxable bonds that pay 10 1/4% interest annually. Given that the risk is the same on the two bonds, which one should Blaze purchase and why
Answer by jim_thompson5910(35256)

Let's assume each bond's face value is $1000


Baptist Health System:
(6 & 1/4)% of 1000 = 0.0625*1000 = 62.50
No taxes paid since these bonds are tax exempt


Community Health Systems:
(10 & 1/4)% of 1000 = 0.1025*1000 = 102.50
33% of 102.5 = 0.33*102.5 = 33.825 = 33.83 in taxes paid
102.50 - 33.83 = 68.67


With the Baptist Health System bond, he nets $62.50 per year in coupon payments.
With the Community Health Systems bond, he nets $68.67 per year in coupon payments.


Both have the same amount of risk, so Blaze should choose the Community Health Systems bond.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------